### v2

# A Spectroscopic Search for Dormant Black Holes in Low-Metallicity Binaries

Abstract
The discovery of the massive black hole (BH) system Gaia BH3 in pre-release Gaia DR4 data suggests that wide BH binaries with luminous companions may be significantly overrepresented at low metallicities. Motivated by this finding, we have initiated a spectroscopic survey of low-metallicity stars exhibiting elevated RUWE values in Gaia DR3, using the FEROS and APF spectrographs. We identify promising BH binary candidates as objects with instantaneously measured radial velocities (RVs) that are very different from their mean RVs reported in Gaia DR3. Thus far, we have observed over 500 targets, including a nearly complete sample of stars with , RUWE , and . Our search has yielded one promising target exhibiting slow acceleration and an RV more than 98 km s-1 different from its DR3 mean RV, as well as dozens of other candidates with smaller RV discrepancies. We quantify the sensitivity of our search using simulations, demonstrating that it recovers at least half of the BH companions within our selection criteria. We make all the spectra and RVs from our survey publicly available and encourage further follow-up.
show]pnagaraj@caltech.edu
1 Introduction
The Milky Way (MW) is thought to contain stellar-mass black holes (BHs; e.g., A. Olejak et al., 2020). Only a tiny fraction of this population has been discovered, as BH searches are severely affected by observational incompleteness. For five decades, the only known Galactic BHs were in close binaries that shine brightly in X-rays due to accretion (e.g., J. M. Corral-Santana et al., 2016). However, these X-ray binaries are a rare outcome of binary evolution, and are expected to be vastly outnumbered by both wider binaries hosting non-accreting (“dormant”) BHs (e.g., S. F. Portegies Zwart et al., 1997; C. Chawla et al., 2022) and isolated BHs (e.g., C. Y. Lam et al., 2022; K. C. Sahu et al., 2022).
The Gaia mission (Gaia Collaboration et al., 2016) has opened a new window on the population of rare binaries in the Milky Way, with the mission’s 3rd data release (“DR3”, Gaia Collaboration et al. 2023a) increasing the number of well-characterized binary orbits by more than an order of magnitude (Gaia Collaboration et al., 2023b). Gaia produces catalogs of photometrically, spectroscopically, and astrometrically variable binaries (Gaia Collaboration et al., 2023b; N. Mowlavi et al., 2023; R. Gomel et al., 2023; L. Eyer et al., 2023; J.-L. Halbwachs et al., 2023), and several recent studies have searched for dormant BHs in these catalogs (e.g., K. El-Badry & H.-W. Rix, 2022; S. Shahaf et al., 2023; P. Nagarajan et al., 2023). In contrast to most other methods, Gaia astrometry is most sensitive to binaries with relatively wide orbits (e.g., J.-L. Halbwachs et al., 2023; K. El-Badry, 2024). Notably, DR3 astrometry has led to the discovery and characterization of two nearby dormant BHs (“Gaia BHs”, K. El-Badry et al. 2023a; S. Chakrabarti et al. 2023; P. Nagarajan et al. 2024b; A. Tanikawa et al. 2023; K. El-Badry et al. 2023b) and a population of dormant neutron stars (NSs) (“Gaia NSs”, K. El-Badry et al. 2024a, b) in au-scale orbits. Population synthesis studies calibrated to the DR3 detections predict discovery of dozens more Gaia BHs in future data releases (e.g., P. Nagarajan et al., 2025).
Recently, Gaia Collaboration et al. (2024) announced the discovery of Gaia BH3, a BH orbited by a red giant in an yr orbit, in pre-release astrometry from the mission’s 4th data release (“DR4”). Gaia BH3 hosts the most massive known stellar-mass BH in the Milky Way, and the lowest-metallicity (, see also E. Balbinot et al. 2024) luminous companion to a BH, providing empirical evidence that low-metallicity massive stars leave behind massive BHs. Because Gaia BH3 is in the typical mass range of the merging BHs detected via gravitational waves (GWs) at extragalactic distances (e.g., R. Abbott et al., 2023), the system also provides tantalizing evidence that low-metallicity massive stars are the progenitors of GW events.
Since Gaia BH3 is relatively bright (), there may be many similar, but fainter, systems waiting to be discovered. Indeed, for a Kroupa initial mass function (IMF) (P. Kroupa, 2001), there are times more low-mass metal-poor main sequence stars in our solar neighborhood than there are red giants similar to the companion star in Gaia BH3. Hence, even within 590 pc (the distance to Gaia BH3), there likely exist other massive BHs that are orbited by metal-poor stars still on the main sequence. At the distance and extinction of Gaia BH3, such companions would have for , readily accessible to Gaia astrometry and spectroscopic follow-up.
[…]
Previous studies have exhaustively searched the catalog of DR3 orbital solutions for dormant BHs and NSs in wide orbits (J. J. Andrews et al., 2022; S. Shahaf et al., 2023; K. El-Badry et al., 2023a, b, 2024a, 2024b). In this work, we perform spectroscopic follow-up on a sample of low-metallicity stars with acceleration solutions or high astrometric excess noise (Renormalized Unit Weight Error, or “RUWE”) values in Gaia DR3, indicating binarity. This investigation cannot be carried out with Gaia data alone. Gaia DR3 includes mean radial velocities (RVs) for stars with , but these represent an average RV measured from spectra coadded over several years (D. Katz et al., 2023). Epoch-level RVs are only measured for stars with , and are not included in the DR3 catalog (D. Katz et al., 2023). External, time-resolved RVs are thus crucial for constraining the possible nature of the companions.
In Section 2, we describe and motivate our search method. In Section 3, we detail our sample selection process. In Section 4, we describe the spectra obtained from our follow-up campaign. In Section 5, we analyze the collected RVs and identify promising compact object candidates for further follow-up. We discuss the completeness and expected yields of our search in Section 6. Finally, we summarize our results in Section 7.
2 Search Strategy
We design our search to be sensitive to Gaia BH3-like binaries. We use the gaiamock code (K. El-Badry et al., 2024), which forward-models Gaia epoch astrometry and model fitting, to simulate epoch astrometry and a radial velocity (RV) curve for Gaia BH3 as it would have been observed in Gaia DR3. We show the epoch astrometry and RV curve in Figure 1, with times provided relative to the DR3 reference epoch of Julian Date (JD) . The epoch astrometry was insufficiently constraining for Gaia BH3 to receive an orbital solution in DR3. However, the source had RUWE , indicative of a poor single-star astrometric fit (L. Lindegren, 2018). In addition, the star’s RV is predicted to be different from the reported mean RV in DR3 by km s-1 over of its orbit. This suggests that a search strategy based on measuring epoch RVs for low-metallicity stars with high RUWE values in DR3, and comparing those RVs to the corresponding mean RVs reported in DR3, could lead to the discovery of BH or NS binaries that did not receive orbital solutions in DR3. Indeed, RV follow-up complements and validates RUWE-based filtering, with even a few RVs being able to break the degeneracy between low companion mass, long-period systems and high companion mass, short-period systems that both have elevated RUWE. The interpretation of any observed RV discrepancy between a binary’s epoch RVs and its mean DR3 RV depends on its true orbital period (see Section 6.3).
To explore what types of BH binaries can be detected via elevated RUWE, we investigate the variation of RUWE with orbital period, distance, and BH mass. We randomly sample sky positions from our catalog of candidate low-metallicity binaries (see Section 3), orbital eccentricities from a thermal distribution, and isotropic binary orientations. We adopt default values of d, kpc, and , typical of binaries that our search strategy is sensitive to, and predict the RUWE in DR3 using gaiamock. We vary one quantity at a time in each panel in Figure 2. We plot curves separately for dwarfs and giants, assuming absolute magnitudes of for dwarfs and for giants. Initially, RUWE increases as the orbital period and projected semi-major axis of the photocenter increase, peaking at days. At d, however, DR3’s limited observing baseline causes RUWE to decrease with orbital period instead. RUWE also decreases as the distance to a binary increases, since this causes the projected semi-major axis of the photocenter to decrease. RUWE is only weakly sensitive to BH mass.
In general, giants lead to slightly larger RUWE values than dwarfs, though there is a somewhat non-intuitive dependence of RUWE on distance in Figure 2. This behavior is determined by how the epoch astrometry uncertainties vary with apparent -band magnitude. The uncertainties are smallest at –, and rise at brighter or fainter magnitudes outside this range (B. Holl et al., 2023). As sources become too bright, uncertainties rise and RUWE falls, causing giants to have smaller RUWE values than dwarfs at close distances. Companions to giants are more detectable than companions to dwarfs only at large distances ( kpc), because the uncertainties for dwarfs increase rapidly at those distances.
Our main sample includes a selection cut of RUWE (see Section 3 for details). This implies that we are sensitive to orbital periods of – d and distances out to kpc. That being said, binaries that have elevated RUWE can show no detectable RV discrepancy if they have high eccentricities (depending on DR3 phase coverage) or face-on orbits. We quantify this in Section 6.
3 Sample Selection
We select candidate metal-poor binaries using two sets of simple cuts on RUWE and the metallicity. We adopt the data-driven metallicities derived by R. Andrae et al. (2023), who use the XGBoost algorithm (T. Chen & C. Guestrin, 2016) to infer stellar parameters from low-resolution Gaia XP spectra. These metallicities are reliable because R. Andrae et al. (2023) train their model on externally derived, high-fidelity stellar parameters from APOGEE (Abdurro’uf et al., 2022) and augment their training data with a set of very-metal-poor stars that have values similar to those in our sample. R. Andrae et al. (2023) also use cross-validation to demonstrate the reliability of their model predictions.
3.1 sample
We selected all sources from the R. Andrae et al. (2023) catalog with , , and the following quality cuts:
Apparent magnitude , to allow for RV measurement with small telescopes.
parallax mas, corresponding to a distance limit of kpc.
, to exclude marginally resolved wide binaries with elevated RUWE due to blending rather than orbital motion (L. Lindegren, 2022).
Color , to exclude hot and cool stars for which the metallicities from R. Andrae et al. (2023) are unlikely to be reliable.
, where is the source’s tangential velocity, and is the source’s total proper motion. This excludes sources with disk-like kinematics, which are more likely to be metal-rich contaminants.
A measured radial velocity in Gaia DR3, so that an RV change can be measured.
, excluding sources with small RV shifts measured during the DR3 observing baseline. This cut applies only to the minority of sources with , for which epoch RVs were measured in DR3.
These cuts yielded a total of 595 sources. Of these, 139 already have an orbital solution reported in the gaiadr3.nss_two_body_orbit catalog. Most of these are astrometric Orbital solutions and have been investigated elsewhere (e.g., A. Tanikawa et al., 2023; K. El-Badry et al., 2023b). None imply massive dark companions.
We measured at least one RV for 425 of the remaining 456 candidates. We have thus observed % of the remaining 456 objects in the sample, or % of 595 when we include the objects with orbital solutions. Of the 456 candidates lacking orbital solutions, 108 have acceleration or variable acceleration solutions reported in the gaiadr3.nss_acceleration_astro catalog. 102 of these sources are in our observed sample. The acceleration solutions suggest that these are long-period ( d; K. El-Badry et al. 2024) binaries, but are insufficient to constrain the companion masses on their own.
3.2 , large sample
Once most of the sample had been observed, we selected additional metal-poor candidates with a selection designed to prioritize large photocenter orbits. We began by selecting sources with according to R. Andrae et al. (2023) and , and applied the same cuts on , , , ipd_frac_multi_peak, DR3 RV, color, and rv_amplitude_robust that are used in the sample. We removed any overlapping sources that were already present in the sample.
Motivated by K. El-Badry et al. (2024) (their Figure 8), we predict the approximate photocenter orbit semi-major axis of each binary from RUWE as
where is the along-scan measurement uncertainty per CCD (e.g. Figure 1 of K. El-Badry et al. 2024) and is a RUWE-informed estimate of the photocenter semi-major axis. We then selected targets satisfying
These cuts yielded a total of 292 sources, of which 31 already had an orbital solution reported in the gaiadr3.nss_two_body_orbit catalog. As before, these sources have already been investigated elsewhere; none harbor massive dark companions.
Luminous companions yield smaller photocenter orbits at fixed mass than dark companions, and so this selection is designed to target binaries with large photocenter orbits corresponding to dark and massive companions. Equation 1 will underestimate in cases where the period significantly exceeds the 1000-d DR3 observing baseline, so we expect most binaries selected in this way to have periods of order 1000 d. Given that the binaries’ periods are not known a priori and there is significant scatter in the relation between RUWE and (C. Y. Lam et al., 2025; K. El-Badry, 2025), it is inevitable that some luminous binaries with smaller orbits will enter the sample as well.
We measured at least one RV for 103 of the remaining 261 candidates. We have thus observed % of the remaining 261 objects in the , large sample, or % of 292 when we include the objects with orbital solutions. Of the 261 candidates lacking orbital solutions, 4 have acceleration or variable acceleration solutions reported in the gaiadr3.nss_acceleration_astro catalog. 2 of these sources are in our observed sample.
3.3 Summary of the sample
Basic properties of our sample of low-metallicity stars with elevated RUWE are summarized in Figure 3. We exclude candidates with orbital solutions. In the upper left panel, we show an extinction-corrected color-magnitude diagram, with a random comparison sample of stars within 100 pc for reference. The extinctions are calculated based on reddenings retrieved from the 3D dust map of G. M. Green et al. (2019). As expected, the low-metallicity stars are bluer than the comparison sample. In the upper right panel, we plot distance versus apparent magnitude for all targets. Our targets tend to lie between – and have typical distances of – kpc. In the lower left panel, we plot metallicity versus RUWE. All observed targets have both [M/H] and high RUWE values indicating binarity. The structure in this panel arises because we use different sample selection criteria for binaries with and . Finally, in the lower right, we show the sky distribution of all targets. The targets are distributed approximately randomly, as expected for halo stars. Some structure is evident, perhaps due in part to our cut of km s-1, which excludes stars on disk-like orbits.
Our selection cuts depend on the availability of robustly measured parallaxes, but DR3 parallax uncertainties for sources with elevated RUWE are underestimated by a factor of – (K. El-Badry, 2025). After inflating the parallax uncertainties of our targets based on the prescription of K. El-Badry (2025), we find that still have parallax to parallax uncertainty ratios , implying that most of our candidates still have reasonably well-constraned distances.
In summary, we obtain at least one epoch RV for a majority of low-metallicity stars with and RUWE . We also obtain epoch RVs for an additional subset of sources with , RUWE , and large photocenter orbits. Most of the binaries in our observed sample are expected to have orbital periods days, though some are likely shorter-period binaries or higher-order multiples that failed one or more of the quality cuts imposed on DR3 astrometric solutions. DR3 reports mean RVs for stars with and RV variability information for stars with . For the typical colors of targets in our sample, these correspond roughly to and , respectively. A complimentary sample of low-metallicity stars with potential BH companions will be investigated in Lam et al. (forthcoming).
4 Data
4.1 Spectroscopic Follow-up
We obtained 657 spectra for 528 targets, years after the median epoch of the time series data in Gaia DR3. Of these spectra, 499 were obtained using FEROS, and 158 were obtained using APF. We used a typical exposure time of 1800s for targets with , and shorter and longer exposure times for brighter and fainter objects. We make all our spectra publicly available at https://doi.org/10.5281/zenodo.16888415.
4.1.1 FEROS
We used the Fiberfed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) on the 2.2m ESO/MPG telescope at La Silla Observatory for targets at (programs 113.26XB.001, 114.27SS.001, and 115.28KE.001). The resulting spectra have resolution over a spectral range of – nm, and a typical SNR of 4 at 5178 Å. We reduced the data using the CERES pipeline (R. Brahm et al., 2017). The pipeline performs bias subtraction, flat fielding, wavelength calibration, and optimal extraction, and measures and corrects for small shifts in the wavelength solution over the course of a night using simultaneous observations of a ThAr lamp with a second fiber. For each spectrum, we estimated the continuum using a median filter, and then performed continuum normalization.
4.1.2 APF
We used the Levy spectrometer on the 2.4m Automated Planet Finder (APF; M. V. Radovan et al., 2010; S. S. Vogt et al., 2014) telescope at Lick Observatory for targets at . We used a 2” slit, yielding spectra with resolution over a wavelength range of – nm and a typical SNR of 8 at 5178 Å. We reduced the data using the California Planet Search pipeline (A. W. Howard et al., 2010; B. J. Fulton et al., 2015). We measured and corrected for small shifts in the wavelength solution over the course of a night using telluric absorption lines (see P. Nagarajan et al. 2023 for a description of the method). As with the FEROS spectra, we estimated the continuum of each spectrum using a median filter, and then performed continuum normalization.
4.2 Radial Velocities
We measured RVs via comparison with a synthetic template spectrum. Specifically, we measured RVs based on two consecutive orders that display prominent absorption features. We used orders 9 and 10 for FEROS and orders 34 and 35 for APF, corresponding to wavelength regions of 5048–5362 Å and 5137–5279 Å, respectively. These wavelength regions include the Mg I b triplet, which in most of our targets is among the strongest features in the optical. We experimented with a variety of spectral regions and found these to be the most robust; the Balmer lines are too broad, and the Mg I lines are among the strongest lines in metal-poor stars. Many other orders do not contain detectable lines in most of our targets.
We calculate the chi-squared statistic as a function of RV shift, defined as:
where is the observed flux, is the template flux, shifted to the appropriate RV, and is the uncertainty in the observed flux at wavelength .
For each target, we chose the template spectrum as the BOSZ Kurucz (R. L. Kurucz, 1979, 1993; S. Mészáros et al., 2024) spectrum with stellar parameters closest to those derived by R. Andrae et al. (2023). The template spectra cover a grid of metallicity spaced by , effective temperature spaced by K, and surface gravity spaced by . The template spectra are at a resolution of and assume no rotational broadening.
The best-fit RV is the shift for which the statistic is minimized. Uncertainties on RVs are derived based on the width of the minimum, i.e., the difference at which the statistic changes by one. All of our RVs are reported in Table 3 in Appendix A.
We manually inspected the template fit in all cases where the epoch RV differed from the mean DR3 RV by more than 10 km s-1 (see Table 1). We do not account for the error in the mean DR3 RV when adopting this threshold, since that error often arises from orbital motion rather than statistical uncertainty. During this vetting process, we visually identified a few objects where a joint fit to both orders produced a less accurate RV than a fit to only one of the orders. In these cases, we adopt the more accurate RV derived from the best fit to one of the orders alone.
We show the observed FEROS spectrum for a typical target, along with the template spectrum used to derive the corresponding best-fit RV, in Figure 4. We plot a spectral region centered on the Mg I b triplet, from which most of the signal used to derive the RV arises. We also show the statistic as a function of RV shift. The location of the minimum in this curve corresponds to the best-fit RV.
4.3 Flagged Spectra
We show the observed FEROS spectrum and statistic as a function of RV shift for a target for which the derived RV is unreliable in Figure 5. The statistic curve clearly has two local minima, suggesting that the system is a double-lined spectroscopic binary (i.e., SB2). Visually, the single-star template shifted into the lab frame does not seem to be a good fit to the observed spectrum for either plausible RV.
Some spectra have low signal-to-noise values (i.e., SNR at 5200 Å) or are likely contaminated with moonlight (i.e. have barycenter correction within km s-1 of the derived RV). Other spectra do not have prominent absorption lines due to high effective temperatures (i.e., the source is a hot star) or rapid rotation (i.e., the source has a high ). In all these cases, we flag the derived RV as unreliable, remove the target from our sample of promising candidates, and report the flag in Table 3.
5 Results
We compare our derived RVs to the mean RVs reported in DR3 in the top panel of Figure 6. We plot good RVs in blue, and RVs for spectra that were flagged in orange. We show a residual plot in the bottom panel. In most cases, the epoch RVs are in good agreement with the mean RVs. In several cases, the two RVs are discrepant beyond the statistical uncertainties, suggesting the possibility of a hidden compact object companion.
[…]
If we account for the uncertainties on our RV measurements and the reported Gaia DR3 RVs, only 49 of the 67 interesting candidates listed in Table 1 have RV discrepancies at the level. However, these uncertainties are dominated by the uncertainty in the Gaia DR3 mean RV, which is often a result of orbital motion rather than the statistical uncertainty on an epoch RV measurement (see Section 6 for further details on the process by which DR3 RVs are computed). As such, we still include the 18 candidates whose largest RV shift is statistically consistent with zero to within , as we do not think that this substantially undermines the evidence indicating the presence of a potential compact object companion.
Of the flagged objects plotted in orange in Figure 6, 52 correspond to low-SNR spectra, while 4 have spectra that are potentially contaminated with moonlight. The remainder were flagged after manual vetting, with 11, 8, and 2 objects determined to be double-lined spectroscopic binaries, hot stars, and fast rotators, respectively. Many of the flagged objects have large RV discrepancies that we find to be spurious, and many of the candidates with large RV discrepancies are flagged. However, several genuine candidates survive the vetting process, and are listed in Table 1.
5.1 Objects with Multiple RVs
We are attempting to obtain at least one additional follow-up RV for each target with RV discrepancy km s-1. We plot RVs over time for two typical targets in Table 1 with RV discrepancy km s-1 in Figure 7. The target in the left panel displays RV variability on short timescales. This suggests that the source is an unresolved hierarchical triple; in this scenario, the observed RV variability arises from the inner binary, while the large RUWE value is due to the motion of the outer orbit. The target in the right panel displays slower RV variability that suggests a long orbital period, consistent with our expectations for typical binaries in our sample.
For one target in our sample, Gaia DR3 2123589958955565824, we measured 11 RVs, which is enough to constrain the full orbital solution. We use ensemble MCMC sampling (emcee; D. Foreman-Mackey et al., 2013) with 64 walkers and 10000 iterations to fit the RVs with a two-body Keplerian orbit and show the best-fit RV curve in Figure 8. We find a short orbital period of d and a moderate eccentricity of . The best-fit RV semi-amplitude of km s-1 leads to a mass function of and a minimum companion mass of , assuming a luminous star mass of . This makes the system unlikely to contain a BH. Using gaiamock, we find that the predicted DR3 RUWE for the derived orbit and typical inclinations is , much smaller than the observed DR3 RUWE of . We conclude that the system is likely a hierarchical triple, with the RVs tracing the motion of the inner binary.
While the total number of RV epochs per target ranges from –, and we observe some RV variability in all targets, we have not yet fully solved the orbit of any of our other targets. For targets with orbital periods days, a few RVs taken over a short time frame are insufficient to fit a full orbital solution and tightly constrain the companion mass. Since extended RV monitoring is necessary to determine the orbital period and confirm or rule out a BH or NS companion, we will continue following up interesting candidates over the next several months. Upon release of DR4, we will derive constraints on the companion masses by jointly fitting our long-baseline measurements with epoch Gaia astrometry.
| Gaia DR3 ID | Most Discrepant FEROS or APF RV | Gaia DR3 Mean RV | Largest RV Shift | |
| km s-1 | km s-1 | km s-1 | ||
| (1) | (2) | (3) | (4) | (5) |
| 6769569470180424704 | 20 | |||
| 6645431961700136576 | 4 | |||
| 6393888749512565760 | 2 | |||
| 2123589958955565824 | 11 | |||
| 1109010106866929536 | 8 | |||
| 6714154840051650560 | 1 | |||
| 4318465066420528000 | 2 | |||
| 5044327218140737536 | 1 | |||
| 4466423837863123328 | 3 | |||
| 6463410835377385472 | 2 | |||
| 5817399066319766912 | 2 | |||
| 3190523740098447744 | 2 | |||
| 3252546886080448384 | 1 | |||
| 1183864824124929792 | 2 | |||
| 5767212132939510400 | 3 | |||
| 5470326985368066304 | 1 | |||
| 2474847090389656064 | 3 | |||
| 4449971364341398272 | 3 | |||
| 6820303307501354496 | 1 | |||
| 2788486767941706624 | 3 | |||
| 677442515834924416 | 3 | |||
| 6177352097769902208 | 2 | |||
| 1589175287211605376 | 1 | |||
| 6822730754297502720 | 2 | |||
| 3529854748080356352 | 1 | |||
| 3195269030422383488 | 1 | |||
| 6402817333686791296 | 3 | |||
| 2402811860245032448 | 3 | |||
| 2429303635139814400 | 1 | |||
| 1294817813795066880 | 1 | |||
| 1694181735744784768 | 1 | |||
| 2619820000613684992 | 1 | |||
| 5510992972680488832 | 1 | |||
| 2438001734187492736 | 2 | |||
| 6689320926830067328 | 1 | |||
| 6646169837081308032 | 1 | |||
| 5889520363370193536 | 1 | |||
| 4324114150961924992 | 2 | |||
| 4036323751356613120 | 2 | |||
| 1445952348525801344 | 1 | |||
| 1363047492460703744 | 1 | |||
| 1028448477581901184 | 1 | |||
| 4404085239419591552 | 2 | |||
| 6657756868572776832 | 2 | |||
| 3838404232257203072 | 1 | |||
| 2281492851046104064 | 1 | |||
| 1805460872350903168 | 1 | |||
| 1911515915671512064 | 1 | |||
| 3619929278952732928 | 1 | |||
| 4792163228561925504 | 1 | |||
| 2835800093315146496 | 2 | |||
| 1343879087778316544 | 1 | |||
| 4344582418904624256 | 1 | |||
| 4480417902167849728 | 3 | |||
| 5000838062927554176 | 1 | |||
| 4463022464283196928 | 1 | |||
| 6288565568615343488 | 1 | |||
| 3638399772924914176 | 1 | |||
| 1558668134509319040 | 2 | |||
| 5059047789053289856 | 1 | |||
| 3270069940330680576 | 2 | |||
| 5186664770791126912 | 1 | |||
| 5300091008802250112 | 1 | |||
| 6903293654893289856 | 1 | |||
| 1115628342232842368 | 10 | |||
| 1698379915953323008 | 1 | |||
| 2292990925172085632 | 1 |
5.2 A Promising Candidate
The target Gaia DR3 6769569470180424704 has the largest discrepancy ( km s-1) between its epoch RV(s) and its Gaia DR3 mean RV in Table 1. We obtained several follow-up spectra to investigate the true nature of this object. We present all of our epoch RVs in Figure 9. We find slow acceleration of km s-1 yr-1, with the RV discrepancy growing over time. Extended RV monitoring is warranted.
We obtained a high signal-to-noise spectrum with the Magellan Inamori Kyocera Echelle (MIKE) spectrograph on the Magellan Clay telescope at Las Campanas Observatory (R. Bernstein et al., 2003) on JD 2460480.7435. We used the 0.7” slit with an exposure time of 2700s, yielding spectral resolution over a spectral range of 333–968 nm. The typical SNR per pixel at 5175 Å is . We reduced the spectrum using the MIKE pipeline within CarPy (D. D. Kelson et al., 2000; D. D. Kelson, 2003). To combine the orders, we perform an inverse variance-weighted average, and normalize the resulting spectrum using a polynomial spline fit to continuum wavelengths.
Using the approach described by H. Reggiani et al. (2022), we jointly fit the SED and high-resolution spectrum to derive atmospheric parameters for this target. We also measure the equivalent widths of atomic absorption lines to infer elemental abundances. We provide our best-fit parameters and abundances in Table 2. The reported uncertainties are statistical, and do not include systematic errors. For comparison, R. Andrae et al. (2023) estimate , K, and . The atmospheric parameters we derive are typical for a low-metallicity halo star (given our selection criteria), and are in reasonable agreement with those provided by R. Andrae et al. (2023). The target’s abundance pattern is unremarkable for a low-metallicity halo star. We verify that non-local thermodynamic equilibrium corrections do not strongly affect any of our derived abundances.
| Parameter | Value | |
| (1) | (2) | (3) |
| K | ||
| km s-1 | ||
| Gyr | ||
| 1 | ||
| 2 | ||
| 7 | ||
| 2 | ||
| 1 | ||
| 12 | ||
| 12 | ||
| 5 | ||
| 27 | ||
| 3 | ||
| 3 | ||
| 74 | ||
| 6 | ||
| 9 | ||
| 13 | ||
| 2 | ||
| 3 | ||
| 1 |
We show the position of Gaia DR3 6769569470180424704 on an extinction-corrected color-magnitude diagram in the top panel of Figure 10. As before, the extinctions are calculated based on reddenings retrieved from the 3D dust map of G. M. Green et al. (2019). We confirm that the target is bluer than the main sequence of a 100 pc random comparison sample, but has typical colors for a metal-poor halo star.
To check for any signatures of a double-lined spectroscopic binary, we follow the approach of P. Nagarajan et al. (2024a) by comparing the high-resolution MIKE spectrum to objects with similar atmospheric parameters observed by the GALAH survey (S. Buder et al., 2021). Motivated by the derived stellar parameters in Table 1, we retrieve all GALAH spectra with K, , and . We then degrade the MIKE spectrum to and shift all spectra into their respective rest frames. We search for the closest matching GALAH spectra in pixel space by minimizing the total least-squares deviation in the wavelength range of . We choose these bounds because virtually all of the signal in the overlapping region between the two spectra arises from the H absorption line. We show the MIKE spectrum along with the closest matching GALAH spectrum in the bottom panel of Figure 10. We do not observe any strong emission lines or other unusual spectral features. The close match rules out the presence of a secondary that makes an appreciable contribution to the total light and has strong H absorption.
Based on the top ten matches in GALAH DR3, we empirically derive stellar parameters of K, , , and km s-1 for Gaia DR3 6769569470180424704. While hot and metal-poor stars are relatively rare in GALAH DR3, these values are in agreement with the best-fit derived parameters listed in Table 2.
To summarize, if the Gaia DR3 RV is correct, then Gaia DR3 6769569470180424704 is very likely to host a dormant BH. However, we do not know if the DR3 RV is accurate, and further monitoring of the acceleration or orbital motion of this target is necessary to determine if this is the case. We check whether plausible orbital solutions that are consistent with the Gaia DR3 RV, Gaia DR3 RUWE, and epoch RVs exist, following the approach of Müller-Horn et al. (forthcoming). We find that only solutions with large companion masses, long orbital periods, and high eccentricities are consistent with the observed acceleration, though we defer further analysis of this to future work.
6 Discussion
6.1 What are the false positives?
False positives in our sample could be targets with spurious Gaia DR3 mean RV measurements, or luminous binaries with primary-to-secondary flux ratios high enough to hide the signatures of the faint secondaries. However, we have removed clear double-lined spectroscopic binaries, hot stars, and fast rotators from consideration. While Gaia’s RV processing pipeline could still potentially fail to recover the true mean RV for the targets in our vetted sample, it is clear from Figure 6 that the DR3 RVs are reliable for the vast majority of targets.
Luminous hierarchical triples are another possible source of false positives (e.g., D. Bashi & A. Tokovinin, 2024). Several objects in Table 1 display RV variability that suggests a short orbital period, with a typical example being the candidate in the left panel of Figure 7. While the large RUWE values for these targets are due to the motion of the outer tertiary, the large RV variability observed on short timescales is due to the orbital motion of the inner binary instead. Indeed, triple systems appear as contaminants whenever a sample of binaries is selected based on both high RUWE and large RV variability (e.g., S. Andrew et al., 2022).
6.2 How many low-metallicity stars in our sample have BH companions?
Metal-poor stars in the solar neighborhood are Gyr old (e.g., K. El-Badry et al., 2018; F. Sestito et al., 2021). Using MIST isochrones (J. Choi et al., 2016), we find that metal-poor giants of this age cover a narrow mass range of –. Assuming a Kroupa IMF (P. Kroupa, 2001), there are times more dwarfs with – compared to these giants. Assuming that Gaia BH3 is the only giant-BH system in our observed sample, and that BH companion probability is independent of luminous star mass in our mass range of interest, how many low-metallicity dwarfs in our sample are likely to have BH companions?
To estimate the answer, we simulate a population of BH + luminous star binaries in the Galactic halo. We draw stars from a Kroupa initial mass function (IMF) (P. Kroupa, 2001) between and . We sample Galactocentric radii from a power-law distribution, assuming a halo density profile (e.g., X.-X. Xue et al. 2015; C. T. Slater et al. 2016; A. J. Deason et al. 2018; studies that adopt a double power-law profile find a shallower inner slope and steeper outer slope). We then calculate heliocentric distances and retain the stars that lie kpc away from the Sun, which we place at a distance of 8.122 kpc away from the Galactic Center (GRAVITY Collaboration et al., 2018) and at a height of 20.8 pc above the Galactic mid-plane (M. Bennett & J. Bovy, 2019). We estimate absolute magnitudes using synthetic photometry from a 12 Gyr MIST isochrone with initial metallicity , and use them to calculate apparent -band magnitudes. We assume zero extinction. We draw orbital periods from a log-uniform distribution between and days (reasonable for Gaia BH3-like systems formed in clusters, see e.g., D. Mar´ın Pina et al. 2024) and eccentricities from a thermal distribution. We assign sky locations by randomly sampling right ascensions and declinations from our catalog of candidate metal-poor binaries. Then, assuming that each luminous star has a BH companion, we randomly sample binary orientations and calculate DR3 RUWE values using the gaiamock code (K. El-Badry et al., 2024). Finally, we apply selection cuts of RUWE and to identify the simulated sources that would make it into our sample. We display the properties of our simulated sample of targets in Figure 11, confirming that brighter and closer binaries tend to have larger predicted RUWE values in DR3.
Based on the location of the Hertzsprung Gap in the color-magnitude diagram in the left panel of Figure 11, we now define giants as stars with absolute -band magnitude and dwarfs as stars with . We find that there are times more dwarf-BH systems than giant-BH systems in the simulated low-metallicity sample; while the IMF predicts that there are more than 8 dwarfs for every giant, most of those dwarfs are too faint to make it into the sample. Thus, if Gaia BH3 is the only giant-BH system in our observed sample, we expect there to be undiscovered dwarf-BH systems in our observed sample as well. As discussed in Section 6.3 below, we expect our search strategy to recover about half of these compact object binaries. If there are 5 genuine BH binaries in our sample, then the estimated false positive rate of our candidate sample is , or about .
These predictions depend upon Gaia BH3 being at a “typical” distance ( kpc) for low-metallicity BH binaries, and they could be overestimates if Gaia BH3 is a statistical anomaly. This history of astronomy has witnessed many cases where event rates and space densities based on the first-discovered object were overestimated (e.g., M. Schmidt, 1963; B. P. Abbott et al., 2017), so additional low-metallicity BH detections are needed to refine our predictions.
6.3 Search sensitivity
To characterize the efficacy of our search, we simulate the fraction of hidden BH companions recovered by our search strategy as a function of orbital period (log-spaced between and days) and BH mass (chosen from or ). We adopt the simulated Galactic population of low-metallicity binaries from Section 6.2. For each binary brighter than , we assign sky locations by randomly sampling right ascensions and declinations from our catalog of candidate metal-poor binaries. Then, we randomly sample an eccentricity from a thermal distribution, assign a random binary orientation and phase, and mock observe the binary at a random time.
Using synthetic photometry from the MIST isochrone and Equations 2 and 3 in Gaia Collaboration et al. (2018), we estimate the apparent magnitude of each binary. Then, we simulate a DR3 mean RV for each binary based on the rv_method_used parameter in DR3. In detail, for bright stars (), we compute the median of epoch RVs measured over a d baseline. This is motivated by the fact that the median rv_nb_transits in our observed sample is . For fainter stars (), we instead compute the simulated DR3 RV as the peak of a total cross-correlation function (CCF) constructed from combining 20 individual epoch CCFs. Here, each individual CCF is represented as a Gaussian with a width of km s-1. This latter approach approximates the process actually carried out in DR3, implying that the RV reported in DR3 for faint stars is closer to the mode than to the mean of the epoch RVs.
Using gaiamock, we compute the fraction of mock observed BH binaries that have both predicted DR3 RUWE and a discrepancy between their epoch RV and mean DR3 RV greater than our adopted threshold of km s-1. To derive the uncertainty on this fraction, we repeat our simulation times. We plot our results in Figure 12. We find that, assuming typical orbital periods ( d) and BH masses (, our search strategy recovers about half of BHs in low-metallicity binaries that lie closer than kpc and have luminous companions brighter than . Some real BH binaries may not be recovered if they have orientations close to face-on or are observed at a time during which their epoch RV is similar to their average RV. The search sensitivity peaks at orbital periods of a few hundred days, with the sensitivity being significantly lower at much shorter or longer orbital periods.
7 Conclusion
The discovery of Gaia BH3 in DR4 pre-release astrometry suggests that luminous companions to Gaia BHs may be dramatically overrepresented at low metallicities. To investigate this hypothesis, we have initiated a radial velocity (RV) survey of metal-poor stars with significant astrometric excess noise (RUWE) values in Gaia DR3 using the FEROS and APF spectrographs. We compared these epoch RVs against the mean RV values reported for these targets in DR3 to find potential black hole (BH) and neutron star (NS) candidates. The first results of our survey are reported here; continued monitoring of promising candidates is ongoing. We summarize our main conclusions below.
Experimental design: In Gaia BH3, the red giant’s RV is different from the mean DR3 RV by at least 20 km s-1 over of its orbit, and by at least 10 km s-1 over of its orbit, suggesting that a search strategy for BHs and NSs based on this discrepancy would be effective (Figure 1). Furthermore, for simulated low-metallicity dwarf-BH and giant-BH binaries, the predicted RUWE in DR3 is significantly above a detectability threshold of for typical periods of – d and distances kpc (Figure 2). Hence, to search for dormant BHs, we measure instantaneous RVs for low-metallicity stars with elevated RUWE values in DR3 and check whether they are highly discrepant with their corresponding mean DR3 RVs.
Summary of the sample: We select a sample of bright, metal-poor sources with and excess astrometric noise in DR3, adopting the robust, data-driven metallicities derived from XP spectra by R. Andrae et al. (2023). We use two different selection cuts (Section 3) for targets with and RUWE and targets with , RUWE , and large estimated , respectively. We apply additional quality cuts to exclude marginally resolved binaries and stars with disk-like orbits. We obtained at least one FEROS or APF spectrum for 528 targets out of a sample of 717 candidates, with the remainder scheduled to be observed in the future. The sample of observed targets has , RUWE , –, and is bluer than the main sequence on a color-magnitude diagram (Figure 3).
RV measurements: We measure epoch RVs by minimizing the chi-squared statistic between the target spectrum and the BOSZ template spectrum with stellar parameters closest to those derived from XP spectra for the target by R. Andrae et al. (2023) (Figure 4). Some spectra have low SNR or are contaminated by moonlight, and others show either a lack of prominent spectral lines or strong evidence of two luminous components (Figure 5). We flag these spectra and remove the corresponding RVs from further consideration. All of our RVs are listed in Table 3.
BH/NS candidates: We compare all of our derived RVs against the mean RVs reported in DR3, finding that most are in good agreement (Figure 6). In several cases, the RVs are discrepant by more than 10 km s-1, indicating the potential for a massive, dark companion. We list all promising candidates in Table 1. We obtained additional follow-up spectra for promising targets, but have not yet fully solved the orbits of any candidates (Figure 7). Some targets show short-timescale RV variability, and are likely to be hierarchical triples. Extended RV monitoring will be vital to determine the true nature of these BH or NS candidates.
A promising BH candidate: We identify Gaia DR3 676956947018042470 as having the largest discrepancy ( km s-1) between its epoch RVs and its mean RV reported in DR3. We obtained several follow-up spectra, including a high-resolution MIKE spectrum, to investigate this promising BH candidate (Figure 9). From the MIKE spectrum, we infer best-fit stellar parameters and abundances for the target, finding that they are typical for a low-metallicity halo star given our selection criteria (Figure 10). We measure a slow acceleration of km s-1 yr-1 away from the mean DR3 RV, but without knowledge of whether the reported mean RV is accurate, further monitoring is necessary to determine the candidate’s true nature.
Simulations: To characterize the effectiveness of our search strategy, we mock-observe a simulated population of BH + luminous star binaries in the Galactic halo, finding that if Gaia BH3 is the only BH + giant binary in our sample, then we expect there to be undiscovered BH + dwarf binaries in our sample as well (Figure 11). We also simulated the fraction of BH companions recovered as a function of orbital period and BH mass, finding that we expect to detect at least half of BH companions with periods of order d (Figure 12).
We will continue to follow-up the interesting objects identified in Table 1. Since these targets have acceleration solutions or high RUWE values rather than orbital solutions in DR3, the orbital periods of these systems are likely to be on the order of – years (K. El-Badry et al., 2024). This renders it practically challenging to cover the full RV curves of these targets with spectroscopic follow-up.
However, we will obtain another epoch RV for these targets following the release of DR4, in which many of these objects are likely to receive an orbital solution. This will provide us with at least two RV measurements separated by a multi-year time baseline. We will fit these RVs together with the epoch astrometry in DR4 to constrain the parameters of the low-metallicity binaries with high implied mass functions and uncover any hidden NSs or BHs.
Appendix A Radial Velocities
We provide a summary of all of our measured radial velocities (RVs) in Table 3. We group RVs by target and sort by largest RV discrepancy with the mean RV reported in DR3. We flag unreliable RVs derived from spectra that have either low signal-to-noise (SNR) values at or moonlight contamination that causes the barycenter correction (BC) to be within km s-1 of the derived RV. We also flag RVs derived from spectra visually identified to be of hot stars (hot), fast rotators (), or spectroscopic binaries (SB2).
| Gaia DR3 ID | JD | RV | Instrument | Gaia DR3 Mean RV | Flags | RV Shift |
| d | km s-1 | km s-1 | km s-1 | |||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| 31530905403904512 | 2460559.9042 | FEROS | ||||
| 31530905403904512 | 2460570.8191 | FEROS | ||||
| 75519548132523136 | 2460569.8113 | FEROS | ||||
| 187869776566408960 | 2460575.8128 | APF | ||||
| 207547530009853952 | 2460574.8033 | APF | ||||
| 249249188555702784 | 2460563.9349 | APF | ||||
| 290511523477577600 | 2460553.9909 | APF | ||||
| 316060462855379840 | 2460566.8837 | APF | ||||
| 316671516442699520 | 2460563.9599 | APF | ||||
| 327354611937489024 | 2460563.9116 | APF | ||||
| 363631486266516992 | 2460568.7686 | APF | ||||
| 374072654842343040 | 2460555.8470 | APF | ||||
| 402635355594132224 | 2460550.9485 | APF | ||||
| 462765992951416064 | 2460562.7899 | APF | BC | |||
| 524686025365395328 | 2460547.9620 | APF | hot | |||
| 583416851271964544 | 2460448.4926 | FEROS | ||||
| 589980763891303040 | 2460448.5106 | FEROS | ||||
| 596344015638431104 | 2460738.6926 | FEROS | SNR | |||
| 614970803468896640 | 2460428.5437 | FEROS | ||||
| 619887407151659648 | 2460450.4587 | FEROS | ||||
| 654976435263115520 | 2460737.6410 | FEROS | SNR | |||
| 659538102848458112 | 2460424.4863 | FEROS | ||||
| 677442515834924416 | 2460587.0107 | APF | ||||
| 677442515834924416 | 2460728.8677 | APF | ||||
| 677442515834924416 | 2460760.7451 | APF | ||||
| 719669191180911488 | 2460587.0275 | APF | ||||
| 859588402366008320 | 2460824.7779 | APF | ||||
| 866004705550987264 | 2460577.9367 | APF | SB2 | |||
| 874790215772404608 | 2460577.9583 | APF | ||||
| 890970938125057280 | 2460580.9244 | APF | ||||
| 900812185667820160 | 2460581.9443 | APF | ||||
| 979611641410430336 | 2460580.9015 | APF | ||||
| 988342141692595328 | 2460577.9785 | APF | ||||
| 1028448477581901184 | 2460586.0190 | APF | ||||
| 1028448477581901184 | 2460717.8399 | APF | BC | |||
| 1028448477581901184 | 2460728.8023 | APF | BC | |||
| 1028448477581901184 | 2460760.7307 | APF | BC | |||
| 1030880524287375360 | 2460585.9656 | APF | ||||
| 1052266579499163392 | 2460581.9223 | APF | ||||
| 1073062536268418688 | 2460581.9032 | APF | ||||
| 1087217545885449600 | 2460573.8435 | APF | BC | |||
| 1093810767362290560 | 2460566.9202 | APF | ||||
| 1095232470257179264 | 2460575.8412 | APF | ||||
| 1109010106866929536 | 2460566.9042 | APF | ||||
| 1109010106866929536 | 2460582.7266 | APF | ||||
| 1109010106866929536 | 2460582.7412 | APF | ||||
| 1109010106866929536 | 2460584.7458 | APF | ||||
| 1109010106866929536 | 2460584.7602 | APF | ||||
| 1109010106866929536 | 2460584.7743 | APF | ||||
| 1109010106866929536 | 2460586.7461 | APF | ||||
| 1109010106866929536 | 2460586.7554 | APF | ||||
| 1115628342232842368 | 2460562.7526 | APF | ||||
| 1115628342232842368 | 2460704.6419 | APF | ||||
| 1115628342232842368 | 2460704.6564 | APF | ||||
| 1115628342232842368 | 2460704.6715 | APF | ||||
| 1115628342232842368 | 2460716.8832 | APF | ||||
| 1115628342232842368 | 2460716.8975 | APF | ||||
| 1115628342232842368 | 2460716.9124 | APF | ||||
| 1115628342232842368 | 2460716.9267 | APF | ||||
| 1115628342232842368 | 2460728.8503 | APF | ||||
| 1115628342232842368 | 2460735.9621 | APF | ||||
| 1126103148991284352 | 2460560.7631 | APF | ||||
| 1141203188852721664 | 2460562.7343 | APF | ||||
| 1149756736482072576 | 2460537.7687 | APF | ||||
| 1161094075752619136 | 2460424.7526 | FEROS | SB2 | |||
| 1161094075752619136 | 2460481.6773 | FEROS | SB2 | |||
| 1177824863156087296 | 2460479.6275 | FEROS | ||||
| 1183864824124929792 | 2460485.6361 | FEROS | ||||
| 1183864824124929792 | 2460738.7903 | FEROS | ||||
| 1186550724873052160 | 2460481.5755 | FEROS | ||||
| 1196101907506720000 | 2460480.6215 | FEROS | ||||
| 1232281475056475392 | 2460478.5826 | FEROS | ||||
| 1236373685536404096 | 2460515.4817 | FEROS | ||||
| 1246628349452501120 | 2460740.7890 | FEROS | SNR | |||
| 1255225155992171392 | 2460813.7746 | APF | ||||
| 1281359310434033664 | 2460806.8174 | APF | ||||
| 1283083482105602560 | 2460813.8110 | APF | ||||
| 1284896817297504896 | 2460801.8186 | APF | ||||
| 1290088711204610560 | 2460825.8357 | APF | ||||
| 1294817813795066880 | 2460846.7540 | APF | ||||
| 1319379563529749120 | 2460817.8264 | APF | ||||
| 1324064475199832448 | 2460582.7102 | APF | ||||
| 1343879087778316544 | 2460535.7197 | APF | ||||
| 1360311082896731136 | 2460573.7517 | APF | ||||
| 1363047492460703744 | 2460542.7441 | APF | ||||
| 1373410286472677888 | 2460819.7913 | APF | ||||
| 1394552589404881152 | 2460813.8255 | APF | ||||
| 1397583290127619840 | 2460583.7263 | APF | ||||
| 1419359598952013952 | 2460543.7390 | APF | ||||
| 1422064500635586688 | 2460531.7299 | APF | ||||
| 1445952348525801344 | 2460815.7786 | APF | ||||
| 1475027838306979200 | 2460828.7687 | APF | ||||
| 1479535389306190976 | 2460473.7443 | APF | BC | |||
| 1479535389306190976 | 2460818.8051 | APF | BC | |||
| 1481189192231609472 | 2460820.8115 | APF | ||||
| 1489161373086670848 | 2460820.7600 | APF | ||||
| 1489688863790805376 | 2460818.7585 | APF | ||||
| 1497188495164830592 | 2460818.7846 | APF | ||||
| 1510085903997844864 | 2460797.8060 | APF | ||||
| 1512383367903955840 | 2460811.7790 | APF | ||||
| 1527759075945701376 | 2460829.7525 | APF | ||||
| 1542487652676277760 | 2460829.7880 | APF | ||||
| 1546566669018411264 | 2460862.7825 | APF | ||||
| 1558668134509319040 | 2460807.7762 | APF | ||||
| 1558668134509319040 | 2460811.7571 | APF | ||||
| 1570143840447017984 | 2460786.7921 | APF | ||||
| 1574040337856735616 | 2460820.7839 | APF | ||||
| 1589175287211605376 | 2460806.8404 | APF | ||||
| 1598550960300560896 | 2460568.7448 | APF | ||||
| 1625301043116465536 | 2460538.7154 | APF | BC | |||
| 1650922176599904000 | 2460486.7830 | APF | ||||
| 1650922176599904000 | 2460533.7698 | APF | ||||
| 1657453619266736512 | 2460804.8068 | APF | ||||
| 1663937336255868032 | 2460786.8118 | APF | ||||
| 1685539544286096384 | 2460561.7649 | APF | ||||
| 1694181735744784768 | 2460543.7584 | APF | ||||
| 1698379915953323008 | 2460561.7148 | APF | ||||
| 1711902775103746560 | 2460543.7739 | APF | ||||
| 1723726369097005312 | 2460529.7376 | APF | ||||
| 1731732737892580352 | 2460464.8922 | FEROS | ||||
| 1779469856283608064 | 2460509.7781 | FEROS | ||||
| 1788672665525489920 | 2460622.5354 | FEROS | ||||
| 1805460872350903168 | 2460480.8230 | FEROS | ||||
| 1810297692736381184 | 2460460.8604 | FEROS | ||||
| 1840445472621264896 | 2460535.7920 | APF | ||||
| 1843327159458210816 | 2460525.8416 | APF | ||||
| 1880593731688347264 | 2460526.8949 | APF | ||||
| 1908577986244357632 | 2460548.7804 | APF | ||||
| 1911515915671512064 | 2460551.8352 | APF | ||||
| 1919029703259598336 | 2460555.8084 | APF | ||||
| 1940380874743387776 | 2460544.8901 | APF | ||||
| 1974994776399961216 | 2460541.7816 | APF | ||||
| 1976290172899067264 | 2460526.8769 | APF | ||||
| 1985939452581432320 | 2460541.8157 | APF | ||||
| 2012367039189084928 | 2460498.7167 | APF | ||||
| 2012367039189084928 | 2460541.0222 | APF | ||||
| 2012367039189084928 | 2460541.7998 | APF | ||||
| 2038568298219233152 | 2460528.7766 | APF | ||||
| 2046121840007606144 | 2460531.7873 | APF | ||||
| 2047697989948528640 | 2460542.7809 | APF | ||||
| 2054522692957346944 | 2460529.7828 | APF | ||||
| 2074915468249960448 | 2460536.7596 | APF | ||||
| 2075888368550754048 | 2460531.7708 | APF | ||||
| 2096216586192095360 | 2460548.7354 | APF | ||||
| 2107991153134598144 | 2460561.7451 | APF | ||||
| 2118002687540997504 | 2460537.7302 | APF | ||||
| 2119880309803249664 | 2460532.7342 | APF | ||||
| 2123589958955565824 | 2460535.7718 | APF | ||||
| 2123589958955565824 | 2460587.6801 | APF | ||||
| 2123589958955565824 | 2460733.9132 | APF | ||||
| 2123589958955565824 | 2460788.9790 | APF | ||||
| 2123589958955565824 | 2460801.9420 | APF | ||||
| 2123589958955565824 | 2460828.9601 | APF | ||||
| 2123589958955565824 | 2460843.8971 | APF | ||||
| 2123589958955565824 | 2460854.7649 | APF | ||||
| 2123589958955565824 | 2460865.8171 | APF | ||||
| 2123589958955565824 | 2460876.8165 | APF | ||||
| 2123589958955565824 | 2460886.8202 | APF | ||||
| 2178042207182395008 | 2460533.7984 | APF | ||||
| 2230058934051637760 | 2460526.9145 | APF | ||||
| 2236448775937348352 | 2460525.8244 | APF | ||||
| 2256036476802548608 | 2460529.7530 | APF | ||||
| 2281492851046104064 | 2460527.9022 | APF | ||||
| 2292990925172085632 | 2460528.7569 | APF | ||||
| 2306397751485602944 | 2460517.8670 | FEROS | ||||
| 2306554049640181760 | 2460517.8837 | FEROS | ||||
| 2315687662667389184 | 2460600.6788 | FEROS | ||||
| 2332061310654863360 | 2460479.9237 | FEROS | ||||
| 2344504075725183104 | 2460519.8773 | FEROS | ||||
| 2379115460522368640 | 2460623.6257 | FEROS | ||||
| 2400926678839820800 | 2460517.7812 | FEROS | ||||
| 2402811860245032448 | 2460481.9391 | FEROS | ||||
| 2402811860245032448 | 2460629.5739 | FEROS | ||||
| 2402811860245032448 | 2460633.5548 | FEROS | ||||
| 2409013453718653824 | 2460624.6115 | FEROS | ||||
| 2410710034520071808 | 2460517.7977 | FEROS | ||||
| 2429303635139814400 | 2460629.6413 | FEROS | ||||
| 2438001734187492736 | 2460519.8280 | FEROS | ||||
| 2438001734187492736 | 2460560.6855 | FEROS | ||||
| 2438414532084246528 | 2460518.8925 | FEROS | ||||
| 2474847090389656064 | 2460568.8389 | FEROS | ||||
| 2474847090389656064 | 2460574.7516 | FEROS | ||||
| 2474847090389656064 | 2460597.6945 | FEROS | ||||
| 2482621256072699904 | 2460575.8227 | FEROS | ||||
| 2487686225105416320 | 2460621.5824 | FEROS | ||||
| 2492878703127223424 | 2460569.7471 | FEROS | ||||
| 2511840193465032448 | 2460598.7108 | FEROS | ||||
| 2524030925559608448 | 2460519.8610 | FEROS | ||||
| 2554217295745049856 | 2460519.8444 | FEROS | ||||
| 2566537942009510528 | 2460568.8061 | FEROS | ||||
| 2570244979822205696 | 2460568.8228 | FEROS | ||||
| 2589877859447847424 | 2460580.7619 | FEROS | ||||
| 2613006872517192960 | 2460517.7564 | FEROS | ||||
| 2613006872517192960 | 2460577.7052 | FEROS | ||||
| 2613006872517192960 | 2460577.7682 | FEROS | ||||
| 2618223166132715776 | 2460627.5910 | FEROS | ||||
| 2619117515762724352 | 2460628.5762 | FEROS | ||||
| 2619820000613684992 | 2460561.7327 | FEROS | SNR | |||
| 2619820000613684992 | 2460628.5515 | FEROS | ||||
| 2643383531229432320 | 2460510.8555 | FEROS | ||||
| 2645540051488656768 | 2460624.5877 | FEROS | SNR | |||
| 2667960605407262080 | 2460518.7833 | FEROS | ||||
| 2669312489313259008 | 2460561.6874 | FEROS | SNR | |||
| 2675008681099778944 | 2460510.7357 | FEROS | ||||
| 2678479564430710656 | 2460623.5810 | FEROS | ||||
| 2701466465622378496 | 2460481.8770 | FEROS | ||||
| 2748452621862730240 | 2460629.6198 | FEROS | ||||
| 2754870849189836800 | 2460560.7994 | FEROS | ||||
| 2758595822787228928 | 2460627.5523 | FEROS | ||||
| 2760474269683840256 | 2460560.7021 | FEROS | ||||
| 2765690181047678336 | 2460519.7906 | FEROS | ||||
| 2765690181047678336 | 2460568.7628 | FEROS | ||||
| 2767395106968307712 | 2460624.6350 | FEROS | ||||
| 2781500952095012096 | 2460560.8361 | FEROS | ||||
| 2781500952095012096 | 2460567.7614 | FEROS | ||||
| 2782781024148048000 | 2460579.6568 | FEROS | ||||
| 2783612048780246016 | 2460575.7886 | FEROS | SNR | |||
| 2788486767941706624 | 2460575.8055 | FEROS | ||||
| 2788486767941706624 | 2460580.6963 | FEROS | SNR | |||
| 2788486767941706624 | 2460596.7702 | FEROS | ||||
| 2788486767941706624 | 2460597.6688 | FEROS | ||||
| 2801479903005558528 | 2460555.8271 | APF | ||||
| 2801532198527486976 | 2460550.9327 | APF | ||||
| 2828393130155180800 | 2460629.5372 | FEROS | ||||
| 2835800093315146496 | 2460527.9174 | APF | ||||
| 2835800093315146496 | 2460544.8193 | APF | ||||
| 2843092157510956032 | 2460544.8543 | APF | ||||
| 2843092157510956032 | 2460548.7996 | APF | ||||
| 2865540508457717376 | 2460544.8737 | APF | ||||
| 2892062618706914560 | 2460625.7524 | FEROS | ||||
| 2896172215213785344 | 2460622.7905 | FEROS | SNR | |||
| 2896660398377207040 | 2460622.7726 | FEROS | ||||
| 2899431683075290240 | 2460623.8294 | FEROS | ||||
| 2905430824673932672 | 2460578.8818 | FEROS | ||||
| 2912512023155224064 | 2460632.6903 | FEROS | ||||
| 2949128990178535552 | 2460712.7325 | FEROS | SNR | |||
| 2950822543021900288 | 2460712.7081 | FEROS | ||||
| 2967475612996185728 | 2460632.6013 | FEROS | SNR | |||
| 2972713033554244480 | 2460623.7808 | FEROS | ||||
| 3008540860462483456 | 2460632.6359 | FEROS | SNR | |||
| 3035481399307040384 | 2460428.5127 | FEROS | ||||
| 3048420147209369600 | 2460732.6603 | FEROS | ||||
| 3078328340945181952 | 2460428.4890 | FEROS | ||||
| 3107577721076953216 | 2460735.6847 | FEROS | SNR | |||
| 3114270895026030720 | 2460734.5719 | FEROS | SNR | |||
| 3146951129379391232 | 2460735.6428 | FEROS | SNR | |||
| 3150574398150536832 | 2460734.6247 | FEROS | SNR | |||
| 3185179705568457344 | 2460626.7658 | FEROS | SB2 | |||
| 3190523740098447744 | 2460572.8677 | FEROS | ||||
| 3190523740098447744 | 2460574.8720 | FEROS | ||||
| 3192689056450786304 | 2460572.8847 | FEROS | SB2 | |||
| 3195269030422383488 | 2460622.7120 | FEROS | ||||
| 3195269030422383488 | 2460633.5916 | FEROS | SNR | |||
| 3205847878470249088 | 2460624.6800 | FEROS | ||||
| 3207170625318358784 | 2460424.4686 | FEROS | ||||
| 3214683828787855360 | 2460624.6987 | FEROS | ||||
| 3239331546588730752 | 2460630.6720 | FEROS | ||||
| 3246885466349457536 | 2460572.8575 | FEROS | ||||
| 3252546886080448384 | 2460622.7369 | FEROS | ||||
| 3258029325934167168 | 2460629.7018 | FEROS | ||||
| 3266994193990116608 | 2460572.8197 | FEROS | ||||
| 3270069940330680576 | 2460559.8769 | FEROS | ||||
| 3270069940330680576 | 2460580.8054 | FEROS | ||||
| 3289282909594133120 | 2460630.6899 | FEROS | ||||
| 3326250307290068224 | 2460712.6276 | FEROS | ||||
| 3334451190860930816 | 2460632.7085 | FEROS | ||||
| 3383997379534709504 | 2460579.9028 | APF | ||||
| 3449708077060728832 | 2460579.0137 | APF | ||||
| 3460512226209856128 | 2460449.6214 | FEROS | ||||
| 3463596081447832704 | 2460740.6636 | FEROS | SNR | |||
| 3463596081447832704 | 2460740.6689 | FEROS | SNR | |||
| 3463780112206037120 | 2460449.5844 | FEROS | ||||
| 3485639846475600000 | 2460733.7985 | FEROS | ||||
| 3524477002349225984 | 2460740.6931 | FEROS | SNR | |||
| 3529854748080356352 | 2460509.4983 | FEROS | ||||
| 3577444188310927744 | 2460427.6929 | FEROS | ||||
| 3584915266743410048 | 2460459.5458 | FEROS | ||||
| 3585663759283867008 | 2460479.5475 | FEROS | ||||
| 3594336603925116672 | 2460739.7846 | FEROS | SNR | |||
| 3618015033504305408 | 2460733.8281 | FEROS | ||||
| 3618107353825818880 | 2460425.7050 | FEROS | ||||
| 3618107353825818880 | 2460509.5217 | FEROS | ||||
| 3619929278952732928 | 2460481.6360 | FEROS | ||||
| 3636369696502927104 | 2460739.8331 | FEROS | ||||
| 3638399772924914176 | 2460480.5636 | FEROS | ||||
| 3651926239847699968 | 2460425.5988 | FEROS | ||||
| 3651926239847699968 | 2460509.5888 | FEROS | ||||
| 3655969212822303616 | 2460428.7253 | FEROS | ||||
| 3657930157450714752 | 2460428.6729 | FEROS | ||||
| 3679906268154294016 | 2460480.5422 | FEROS | ||||
| 3690660213428217344 | 2460462.5791 | FEROS | ||||
| 3718625668909565952 | 2460427.6692 | FEROS | ||||
| 3720871700582825088 | 2460485.6118 | FEROS | ||||
| 3736156767635254272 | 2460480.5043 | FEROS | ||||
| 3737235663419837568 | 2460462.5552 | FEROS | ||||
| 3737235663419837568 | 2460480.4853 | FEROS | ||||
| 3741881684162720768 | 2460481.4914 | FEROS | ||||
| 3816735508068512512 | 2460738.7418 | FEROS | ||||
| 3824738543969555840 | 2460449.4811 | FEROS | ||||
| 3832884821324406400 | 2460425.5153 | FEROS | SB2 | |||
| 3832884821324406400 | 2460478.4713 | FEROS | SB2 | |||
| 3838404232257203072 | 2460424.5229 | FEROS | ||||
| 3842494866483989120 | 2460449.4643 | FEROS | ||||
| 3850064076328182656 | 2460449.4980 | FEROS | ||||
| 3900567355515742976 | 2460425.5521 | FEROS | SB2 | |||
| 3900567355515742976 | 2460509.4580 | FEROS | SB2 | |||
| 3903598369771891072 | 2460459.5079 | FEROS | ||||
| 3906590175271460480 | 2460740.7203 | FEROS | SNR | |||
| 3931304688402594176 | 2460481.4648 | FEROS | ||||
| 4036323751356613120 | 2460517.6421 | FEROS | ||||
| 4036323751356613120 | 2460599.4986 | FEROS | ||||
| 4037752218725563392 | 2460738.8551 | FEROS | ||||
| 4078664878917893632 | 2460427.8245 | FEROS | ||||
| 4219851655185139328 | 2460560.6612 | FEROS | ||||
| 4229000828879015936 | 2460509.7532 | FEROS | ||||
| 4242596049637833472 | 2460509.7196 | FEROS | ||||
| 4242596049637833472 | 2460624.5232 | FEROS | ||||
| 4243579356635038336 | 2460428.8829 | FEROS | ||||
| 4297175497614395008 | 2460518.7477 | FEROS | SNR | |||
| 4297175497614395008 | 2460599.5244 | FEROS | ||||
| 4311422866241470080 | 2460515.6762 | FEROS | ||||
| 4318465066420528000 | 2460425.8072 | FEROS | ||||
| 4318465066420528000 | 2460459.8523 | FEROS | ||||
| 4324114150961924992 | 2460462.7388 | FEROS | ||||
| 4324114150961924992 | 2460736.8534 | FEROS | ||||
| 4341979565646134016 | 2460478.7052 | FEROS | ||||
| 4344582418904624256 | 2460479.6968 | FEROS | ||||
| 4399319028313941760 | 2460485.6876 | FEROS | ||||
| 4404085239419591552 | 2460428.7591 | FEROS | ||||
| 4404085239419591552 | 2460736.8124 | FEROS | ||||
| 4406004814924526848 | 2460425.7244 | FEROS | ||||
| 4406004814924526848 | 2460569.4858 | FEROS | ||||
| 4419230084377663744 | 2460736.8702 | FEROS | ||||
| 4449971364341398272 | 2460424.7693 | FEROS | ||||
| 4449971364341398272 | 2460459.7247 | FEROS | ||||
| 4449971364341398272 | 2460739.8754 | FEROS | ||||
| 4458199593966253184 | 2460462.7074 | FEROS | ||||
| 4458378470765084928 | 2460428.7764 | FEROS | ||||
| 4458669910066960384 | 2460424.7862 | FEROS | SB2 | |||
| 4458669910066960384 | 2460481.7005 | FEROS | SB2 | |||
| 4462674224036731392 | 2460428.7930 | FEROS | ||||
| 4463022464283196928 | 2460485.6700 | FEROS | ||||
| 4464923863483932672 | 2460478.6252 | FEROS | ||||
| 4465345320035111424 | 2460485.6531 | FEROS | ||||
| 4466423837863123328 | 2460478.6553 | FEROS | ||||
| 4466423837863123328 | 2460521.5944 | FEROS | ||||
| 4466423837863123328 | 2460560.5017 | FEROS | ||||
| 4466423837863123328 | 2460740.8790 | FEROS | SNR | |||
| 4480417902167849728 | 2460425.7865 | FEROS | ||||
| 4480417902167849728 | 2460513.5853 | FEROS | ||||
| 4480417902167849728 | 2460600.4939 | FEROS | ||||
| 4486716699344464640 | 2460464.7472 | FEROS | ||||
| 4526671921107968640 | 2460509.6333 | FEROS | SB2 | |||
| 4556877979423587456 | 2460536.7213 | APF | ||||
| 4568304825812360064 | 2460573.7711 | APF | ||||
| 4575229962359762176 | 2460575.7399 | APF | ||||
| 4596554062668403328 | 2460568.7210 | APF | ||||
| 4600988839738641664 | 2460574.7754 | APF | ||||
| 4605537828581921792 | 2460560.7393 | APF | ||||
| 4615537233999076608 | 2460623.7010 | FEROS | ||||
| 4618585973584829440 | 2460575.8436 | FEROS | ||||
| 4623987839851960576 | 2460632.5755 | FEROS | ||||
| 4625065361248005248 | 2460630.7150 | FEROS | ||||
| 4630416817482099712 | 2460623.7181 | FEROS | ||||
| 4633036202073128448 | 2460571.8692 | FEROS | SB2 | |||
| 4651827199427540224 | 2460632.5580 | FEROS | ||||
| 4661235445562538112 | 2460625.6956 | FEROS | ||||
| 4674062382519374208 | 2460629.7191 | FEROS | SNR | |||
| 4678849484347880576 | 2460576.8317 | FEROS | ||||
| 4695299002933110272 | 2460627.6790 | FEROS | ||||
| 4696753927398383872 | 2460624.6560 | FEROS | ||||
| 4696981290083296640 | 2460569.8351 | FEROS | ||||
| 4705880840277690624 | 2460627.7031 | FEROS | ||||
| 4710671206281480192 | 2460569.7286 | FEROS | ||||
| 4719110473421285888 | 2460561.7768 | FEROS | ||||
| 4719110473421285888 | 2460574.8098 | FEROS | ||||
| 4721517579251891712 | 2460570.8754 | FEROS | ||||
| 4728502295587116160 | 2460626.6800 | FEROS | ||||
| 4738910650613013120 | 2460569.7929 | FEROS | ||||
| 4739893854526425216 | 2460570.7776 | FEROS | ||||
| 4746302632927184128 | 2460559.8940 | FEROS | ||||
| 4746302632927184128 | 2460569.7828 | FEROS | ||||
| 4751527580901330816 | 2460579.7541 | FEROS | ||||
| 4753774261113868416 | 2460571.8471 | FEROS | ||||
| 4757764182654737792 | 2460625.7319 | FEROS | ||||
| 4762898249119670400 | 2460631.5962 | FEROS | ||||
| 4770375340505228928 | 2460632.6728 | FEROS | ||||
| 4775145228104639232 | 2460578.8424 | FEROS | ||||
| 4777669947320442752 | 2460626.8083 | FEROS | ||||
| 4790283098038091008 | 2460630.6147 | FEROS | ||||
| 4791430335342393600 | 2460626.7068 | FEROS | ||||
| 4792163228561925504 | 2460631.6130 | FEROS | ||||
| 4798980303733591552 | 2460632.6550 | FEROS | ||||
| 4801894284424445056 | 2460623.7628 | FEROS | ||||
| 4803825026843762560 | 2460632.6187 | FEROS | ||||
| 4805795901436022784 | 2460622.8149 | FEROS | ||||
| 4810532769327201792 | 2460630.6463 | FEROS | ||||
| 4814070963385616384 | 2460600.8083 | FEROS | ||||
| 4815626737978833792 | 2460630.5977 | FEROS | ||||
| 4835367335384068352 | 2460598.7637 | FEROS | ||||
| 4841632386998925440 | 2460630.5618 | FEROS | ||||
| 4867178886835317376 | 2460573.8848 | FEROS | ||||
| 4880417797108293632 | 2460578.8239 | FEROS | ||||
| 4880744901817831552 | 2460578.8602 | FEROS | ||||
| 4884385694054284672 | 2460630.5446 | FEROS | ||||
| 4885231081056967168 | 2460622.6869 | FEROS | ||||
| 4902931740794761728 | 2460599.7656 | FEROS | ||||
| 4903248017890484608 | 2460568.6480 | FEROS | ||||
| 4906613008803987456 | 2460560.8557 | FEROS | ||||
| 4912541472422371840 | 2460561.8290 | FEROS | SNR | |||
| 4929840157581555840 | 2460598.6608 | FEROS | ||||
| 4938691501062803072 | 2460570.7608 | FEROS | ||||
| 4950010114277787776 | 2460567.6458 | FEROS | ||||
| 4951630313380205952 | 2460621.6450 | FEROS | ||||
| 4965484503487580032 | 2460571.8857 | FEROS | ||||
| 4967858555250886272 | 2460568.8840 | FEROS | ||||
| 4973715889913402112 | 2460598.6233 | FEROS | ||||
| 4978614355950125696 | 2460517.9003 | FEROS | SB2 | |||
| 4978614355950125696 | 2460518.9095 | FEROS | SB2 | |||
| 4978614355950125696 | 2460520.8754 | FEROS | SB2 | |||
| 4992290493893169664 | 2460560.8173 | FEROS | ||||
| 5000838062927554176 | 2460598.6416 | FEROS | ||||
| 5034075646601072000 | 2460567.8279 | FEROS | SB2 | |||
| 5035314860629134080 | 2460567.8440 | FEROS | ||||
| 5035671519008704256 | 2460600.7171 | FEROS | ||||
| 5039157932941255296 | 2460626.6611 | FEROS | ||||
| 5041854072891578496 | 2460580.7792 | FEROS | SNR | |||
| 5042687811943245184 | 2460578.7452 | FEROS | ||||
| 5044327218140737536 | 2460629.7470 | FEROS | ||||
| 5044327218140737536 | 2460633.5674 | FEROS | SNR | |||
| 5059047789053289856 | 2460571.8031 | FEROS | ||||
| 5082958696621785984 | 2460600.7910 | FEROS | ||||
| 5083871359992456448 | 2460572.8370 | FEROS | ||||
| 5089796971751769856 | 2460600.7431 | FEROS | ||||
| 5094819506507559296 | 2460578.7690 | FEROS | ||||
| 5119215607942562944 | 2460621.6725 | FEROS | ||||
| 5139848390314545792 | 2460628.7160 | FEROS | SNR | |||
| 5139848390314545792 | 2460629.6845 | FEROS | ||||
| 5141607437120133248 | 2460561.8114 | FEROS | ||||
| 5157067189081730560 | 2460570.8016 | FEROS | ||||
| 5158765384790885760 | 2460571.8241 | FEROS | ||||
| 5183256525263276288 | 2460599.8495 | FEROS | ||||
| 5186664770791126912 | 2460572.8019 | FEROS | ||||
| 5188617434724794624 | 2460450.5197 | FEROS | SNR | |||
| 5195206189434712064 | 2460738.6450 | FEROS | SNR | |||
| 5196065870088374656 | 2460738.6683 | FEROS | SNR | |||
| 5197671122705982080 | 2460448.5942 | FEROS | ||||
| 5202922905635808640 | 2460448.5353 | FEROS | ||||
| 5240547605125005696 | 2460428.6183 | FEROS | hot | |||
| 5274125526303605376 | 2460735.7869 | FEROS | SNR | |||
| 5300091008802250112 | 2460449.5324 | FEROS | ||||
| 5321281033104697088 | 2460450.5021 | FEROS | ||||
| 5326941941734897024 | 2460425.4877 | FEROS | hot | |||
| 5327915383187030784 | 2460459.4617 | FEROS | SNR | |||
| 5334791939429604608 | 2460479.5963 | FEROS | SNR | |||
| 5335252737865276032 | 2460427.5913 | FEROS | ||||
| 5336819782417836288 | 2460425.5348 | FEROS | hot | |||
| 5336819782417836288 | 2460509.4754 | FEROS | hot | |||
| 5348155781656154880 | 2460739.7578 | FEROS | ||||
| 5356371259506596480 | 2460428.5749 | FEROS | hot | |||
| 5356567870254195328 | 2460425.4979 | FEROS | hot | |||
| 5356567870254195328 | 2460478.4882 | FEROS | hot | |||
| 5359619839640581376 | 2460479.5292 | FEROS | hot | |||
| 5367299790914345984 | 2460448.5697 | FEROS | SNR | |||
| 5371704297057727232 | 2460449.6010 | FEROS | ||||
| 5376423126081823360 | 2460448.6549 | FEROS | SNR | |||
| 5376693704723955840 | 2460450.5929 | FEROS | SNR | |||
| 5393856089096403840 | 2460427.6155 | FEROS | ||||
| 5393856089096403840 | 2460448.6306 | FEROS | ||||
| 5399604438965438848 | 2460740.6289 | FEROS | SNR | |||
| 5399748818585456768 | 2460741.8163 | FEROS | SNR | |||
| 5400128459336707072 | 2460449.5675 | FEROS | ||||
| 5432677164616070144 | 2460739.7213 | FEROS | ||||
| 5437197703594043648 | 2460448.5527 | FEROS | ||||
| 5470326985368066304 | 2460448.6137 | FEROS | ||||
| 5485223065822432000 | 2460731.6028 | FEROS | ||||
| 5491120949273989248 | 2460732.6335 | FEROS | ||||
| 5509286152674973184 | 2460731.6429 | FEROS | ||||
| 5510992972680488832 | 2460735.7635 | FEROS | ||||
| 5533294878122904448 | 2460450.4765 | FEROS | ||||
| 5558621205053610368 | 2460625.7764 | FEROS | ||||
| 5560933512367258752 | 2460625.7935 | FEROS | ||||
| 5575760254868795136 | 2460633.8033 | FEROS | ||||
| 5580281034364803072 | 2460633.7614 | FEROS | ||||
| 5604520833353953280 | 2460735.6605 | FEROS | SNR | |||
| 5614140632394855680 | 2460735.7389 | FEROS | ||||
| 5688075127125733504 | 2460449.5152 | FEROS | ||||
| 5713661999660822016 | 2460738.5335 | FEROS | ||||
| 5735304576081570048 | 2460738.7175 | FEROS | ||||
| 5761883109315557888 | 2460739.6874 | FEROS | ||||
| 5764114052767463680 | 2460740.5763 | FEROS | SNR | |||
| 5767212132939510400 | 2460425.7434 | FEROS | ||||
| 5767212132939510400 | 2460481.7543 | FEROS | ||||
| 5767212132939510400 | 2460736.8360 | FEROS | ||||
| 5770704628546574848 | 2460481.5334 | FEROS | SNR | |||
| 5772808715845778176 | 2460519.5012 | FEROS | SNR | |||
| 5783922338860010368 | 2460479.5723 | FEROS | ||||
| 5817399066319766912 | 2460485.7703 | FEROS | ||||
| 5817399066319766912 | 2460598.5128 | FEROS | ||||
| 5831887605960848000 | 2460738.8314 | FEROS | ||||
| 5832426023053419648 | 2460478.6800 | FEROS | ||||
| 5842093547727016960 | 2460510.4699 | FEROS | SNR | |||
| 5851089266522263936 | 2460509.5721 | FEROS | ||||
| 5858256226910097536 | 2460448.6798 | FEROS | ||||
| 5889520363370193536 | 2460738.8074 | FEROS | ||||
| 5915536178508998784 | 2460425.7672 | FEROS | ||||
| 5915536178508998784 | 2460510.6934 | FEROS | ||||
| 5916840787735273216 | 2460464.7714 | FEROS | ||||
| 5923359031943461120 | 2460462.8226 | FEROS | ||||
| 5923503450243018752 | 2460428.8116 | FEROS | ||||
| 6039722860077339136 | 2460740.8552 | FEROS | ||||
| 6072111685190490112 | 2460480.5216 | FEROS | hot | |||
| 6082341236436152576 | 2460740.7465 | FEROS | ||||
| 6104480246655766016 | 2460741.8774 | FEROS | SNR | |||
| 6104628783805040384 | 2460518.5807 | FEROS | SNR | |||
| 6104628783805040384 | 2460712.7988 | FEROS | SNR | |||
| 6109390489486849408 | 2460427.7891 | FEROS | ||||
| 6117757983437538432 | 2460510.6131 | FEROS | ||||
| 6118173976785581568 | 2460740.8137 | FEROS | ||||
| 6133984369878480128 | 2460739.8093 | FEROS | ||||
| 6134138267147056768 | 2460733.7462 | FEROS | ||||
| 6140717877864676352 | 2460481.6117 | FEROS | ||||
| 6154038774769241600 | 2460428.6558 | FEROS | ||||
| 6157786696375384064 | 2460428.6359 | FEROS | ||||
| 6165452942416452864 | 2460449.6425 | FEROS | ||||
| 6177352097769902208 | 2460428.7071 | FEROS | ||||
| 6177352097769902208 | 2460712.7814 | FEROS | ||||
| 6188708953373074048 | 2460449.6612 | FEROS | ||||
| 6214736008513502976 | 2460510.6375 | FEROS | ||||
| 6235652499248057728 | 2460462.7896 | FEROS | ||||
| 6275466158880793856 | 2460427.7231 | FEROS | ||||
| 6275708429396071936 | 2460510.5707 | FEROS | ||||
| 6288565568615343488 | 2460739.8496 | FEROS | ||||
| 6289559179234696960 | 2460428.6904 | FEROS | ||||
| 6319928416243214464 | 2460515.6266 | FEROS | SNR | |||
| 6349293863560840576 | 2460519.6773 | FEROS | SNR | |||
| 6368640114249326976 | 2460481.8249 | FEROS | ||||
| 6374547824585473536 | 2460480.7590 | FEROS | ||||
| 6375836074255319936 | 2460479.8875 | FEROS | ||||
| 6393888749512565760 | 2460513.8346 | FEROS | ||||
| 6393888749512565760 | 2460623.6078 | FEROS | ||||
| 6402359833770432512 | 2460510.7112 | FEROS | ||||
| 6402817333686791296 | 2460485.8242 | FEROS | ||||
| 6402817333686791296 | 2460623.5634 | FEROS | ||||
| 6402817333686791296 | 2460633.5385 | FEROS | ||||
| 6403003597828842240 | 2460481.8594 | FEROS | ||||
| 6409038026880037376 | 2460622.6178 | FEROS | ||||
| 6420147904684678144 | 2460479.7509 | FEROS | ||||
| 6420303519939663488 | 2460625.5259 | FEROS | ||||
| 6421461168246828032 | 2460599.6064 | FEROS | ||||
| 6421465531933455360 | 2460428.9004 | FEROS | ||||
| 6422752304132663680 | 2460479.7769 | FEROS | ||||
| 6431108077108143488 | 2460599.6533 | FEROS | ||||
| 6433084105663790080 | 2460464.7950 | FEROS | ||||
| 6445023942945668992 | 2460427.8556 | FEROS | ||||
| 6451215048105801216 | 2460461.9004 | FEROS | ||||
| 6457366957398956928 | 2460481.8071 | FEROS | ||||
| 6463410835377385472 | 2460558.5699 | FEROS | ||||
| 6463410835377385472 | 2460575.5925 | FEROS | ||||
| 6463410835377385472 | 2460622.5698 | FEROS | SNR | |||
| 6469982100983952896 | 2460623.5466 | FEROS | ||||
| 6473015344324423680 | 2460464.8681 | FEROS | ||||
| 6477293200470828288 | 2460481.8421 | FEROS | ||||
| 6484717029965817984 | 2460462.9089 | FEROS | ||||
| 6484717029965817984 | 2460478.8932 | FEROS | ||||
| 6487033842699994496 | 2460599.7248 | FEROS | ||||
| 6492216047160201728 | 2460623.6514 | FEROS | ||||
| 6504574764016144896 | 2460513.8178 | FEROS | ||||
| 6508841693764232704 | 2460510.7589 | FEROS | ||||
| 6520467998436840448 | 2460510.7930 | FEROS | ||||
| 6528915614992166528 | 2460513.8638 | FEROS | SNR | |||
| 6532929794505116800 | 2460622.6419 | FEROS | ||||
| 6534773125749528960 | 2460626.6443 | FEROS | ||||
| 6538531393931377152 | 2460513.8887 | FEROS | ||||
| 6538531393931377152 | 2460577.7344 | FEROS | ||||
| 6538606396945476736 | 2460625.6136 | FEROS | ||||
| 6565245300240135552 | 2460622.5947 | FEROS | ||||
| 6574816067923362048 | 2460515.8217 | FEROS | ||||
| 6597599529397812608 | 2460510.7758 | FEROS | ||||
| 6597676048534977024 | 2460516.8058 | FEROS | ||||
| 6597676048534977024 | 2460577.6767 | FEROS | ||||
| 6597875094499505280 | 2460626.6193 | FEROS | ||||
| 6598930556942930688 | 2460628.6493 | FEROS | ||||
| 6605719555913798656 | 2460629.6017 | FEROS | ||||
| 6623367718266041344 | 2460509.8637 | FEROS | ||||
| 6626455181276398848 | 2460629.5844 | FEROS | SNR | |||
| 6635741587769528960 | 2460463.8002 | FEROS | ||||
| 6636851819634037632 | 2460478.7667 | FEROS | ||||
| 6639833007973763712 | 2460599.6325 | FEROS | ||||
| 6640871668503771520 | 2460428.8454 | FEROS | ||||
| 6645431961700136576 | 2460428.8285 | FEROS | ||||
| 6645431961700136576 | 2460459.8064 | FEROS | ||||
| 6645431961700136576 | 2460464.6628 | FEROS | ||||
| 6645431961700136576 | 2460478.7980 | FEROS | ||||
| 6645857438339795072 | 2460622.5528 | FEROS | ||||
| 6646169837081308032 | 2460623.5227 | FEROS | ||||
| 6647005122321157120 | 2460464.8188 | FEROS | ||||
| 6647408196416540160 | 2460478.7302 | FEROS | ||||
| 6657756868572776832 | 2460425.6664 | FEROS | ||||
| 6657756868572776832 | 2460481.7712 | FEROS | ||||
| 6661890173299322112 | 2460626.5243 | FEROS | SNR | |||
| 6675903066740376064 | 2460460.9081 | FEROS | ||||
| 6685959341824044032 | 2460428.8625 | FEROS | ||||
| 6688217566909990528 | 2460509.6584 | FEROS | ||||
| 6689320926830067328 | 2460463.8430 | FEROS | ||||
| 6692958519311941760 | 2460462.8541 | FEROS | ||||
| 6694977673633913728 | 2460460.8846 | FEROS | ||||
| 6714154840051650560 | 2460627.5117 | FEROS | ||||
| 6714154840051650560 | 2460633.5140 | FEROS | SNR | |||
| 6752999928728671104 | 2460509.7958 | FEROS | ||||
| 6753027107279926144 | 2460463.8675 | FEROS | ||||
| 6754940639769782400 | 2460464.8433 | FEROS | ||||
| 6769569470180424704 | 2460463.8181 | FEROS | ||||
| 6769569470180424704 | 2460464.6870 | FEROS | ||||
| 6769569470180424704 | 2460467.8426 | FEROS | SNR | |||
| 6769569470180424704 | 2460478.8161 | FEROS | ||||
| 6769569470180424704 | 2460479.7246 | FEROS | ||||
| 6769569470180424704 | 2460480.6921 | FEROS | ||||
| 6769569470180424704 | 2460515.7119 | FEROS | ||||
| 6769569470180424704 | 2460517.7305 | FEROS | ||||
| 6769569470180424704 | 2460518.7098 | FEROS | ||||
| 6769569470180424704 | 2460519.6529 | FEROS | ||||
| 6769569470180424704 | 2460532.7545 | APF | ||||
| 6769569470180424704 | 2460555.7446 | APF | ||||
| 6769569470180424704 | 2460559.7342 | APF | ||||
| 6769569470180424704 | 2460559.7582 | APF | ||||
| 6769569470180424704 | 2460570.5459 | FEROS | ||||
| 6769569470180424704 | 2460580.6043 | FEROS | SNR | |||
| 6769569470180424704 | 2460598.5582 | FEROS | ||||
| 6769569470180424704 | 2460789.7973 | FEROS | SNR | |||
| 6769569470180424704 | 2460790.8146 | FEROS | ||||
| 6769569470180424704 | 2460815.8053 | FEROS | ||||
| 6769569470180424704 | 2460843.6725 | FEROS | ||||
| 6769569470180424704 | 2460847.7174 | FEROS | ||||
| 6769569470180424704 | 2460863.6412 | FEROS | ||||
| 6789503822144617216 | 2460480.8009 | FEROS | ||||
| 6790896731578064640 | 2460479.8226 | FEROS | ||||
| 6792552046335955456 | 2460624.5626 | FEROS | SNR | |||
| 6797640173831774976 | 2460478.8590 | FEROS | ||||
| 6801546326327279104 | 2460479.8014 | FEROS | ||||
| 6803040081593000960 | 2460626.5702 | FEROS | ||||
| 6807367999878036736 | 2460627.5702 | FEROS | ||||
| 6812596551329038208 | 2460628.5256 | FEROS | ||||
| 6815475042707068672 | 2460625.5684 | FEROS | ||||
| 6820063064210514816 | 2460509.8212 | FEROS | ||||
| 6820303307501354496 | 2460628.6249 | FEROS | ||||
| 6822730754297502720 | 2460509.8396 | FEROS | ||||
| 6822730754297502720 | 2460628.6005 | FEROS | ||||
| 6823445986611497728 | 2460626.5949 | FEROS | ||||
| 6829836378256275072 | 2460478.9102 | FEROS | ||||
| 6834540436892911488 | 2460625.5513 | FEROS | ||||
| 6835881321387391872 | 2460480.7802 | FEROS | ||||
| 6843620100836059008 | 2460561.6018 | FEROS | ||||
| 6857258511787349632 | 2460478.8757 | FEROS | ||||
| 6864981928139773696 | 2460478.8409 | FEROS | ||||
| 6897453324003046784 | 2460479.8392 | FEROS | ||||
| 6903293654893289856 | 2460462.8852 | FEROS |